# Exploring Accurate 3D Phenotyping in Greenhouse through Neural Radiance Fields

### v1

@@ III-B1 Reconstruction from Neural Radiance Fields
The traditional NeRF is a 3D reconstruction method that is based on ray tracing and uses the MLP as an implicit representation. Input the point position coordinates ($x,y,z$) and the viewing direction ($\theta,\phi$) into the ray equation $r(t)=o+t\mathbf{d}$, where $o$ is the position coordinate, $\mathbf{d}$ is the viewing direction, $t$ is the distance from the sampling point to the origin. In particular, the position coordinates and viewing direction need to be added to the encoding, and the encoding function is
@@
The Encoding is useful for extracting high-dimensional information for feature enhancement and avoiding similarity.[29] Each pixel point radiates a ray and creates MLP (Multi-Layer Perception), which is trained and saves the weighting parameters. NeRF’s MLP has 10 layers, with 256 neurons in each of the first 9 layers and 128 neurons in the 10th layer. Then volume rendering is used to predict the colour values of the pixel points.[30] The volume rendering function is
@@
where $T(t)$ denotes the transmittance function, which is affected by the opacity of the sampling points and the sampling distance
@@
Since the sampled points are discrete, we need to take an integral of them. So we take an approximate integral to calculate the rendered color values.The approximate equation is
@@
Finally, the loss is calculated from the rendered colour values and backpropagated.
@@
Neural Reconstruction by Neus: Neus is a reconstruction method that allows the surface of an object to be reconstructed more closely to the real surface. It takes the density values used by NeRF (Neural Radiance Fields) and replaces them with the point-to-surface distance (Signed Distance Field)[27]. The surface $\mathbf{S}$ of an object is the function equal to zero, the function is $\mathbf{S}=\left\{\mathrm{x}\in\mathbb{R}^{3}|f(\mathbf{x})=0\right\}.$ However, in 3D ray-tracing scenes, the use of the basic logistic density distribution $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}.$ introduces surface biases that affect the quality of the reconstruction. Therefore, Neus introduces an unbiased and occlusion-aware weight function
@@
The Iterative Closest Point (ICP) algorithm iterates over the nearest points to obtain corresponding points. The point cloud data of different orientations and angles are unified under the same coordinate system through the Euler transformation relationship between corresponding points. This enriches the target point cloud data and makes it more complete. In 3D point cloud acquisition, the positional mapping between two sets of points is represented by displacement and rotation transformations totalling six degrees of freedom.
@@ III-D2 Measurement of Phenotypic data
The 3D model reconstructed using NeRF needs to be converted to actual size for final phenotype estimation. To achieve this, scaling markers that are regular in shape and relatively complete must be found and identified. Accurate calibration plates are used as markers for the calibration reconstruction. The scale factor between the estimated and actual lengths of the calibration plate reconstructed point cloud is then used as the corrected scale of the reconstruction. The estimation result is

### v2

@@ III-B1 Reconstruction from Neural Radiance Fields
The traditional NeRF is a 3D reconstruction method that is based on ray tracing and uses the MLP as an implicit representation. Input the point position coordinates ($x,y,z$) and the viewing direction ($\theta,\phi$) into the ray equation $r(t)=o+t\mathbf{d}$, where $o$ is the position coordinate, $\mathbf{d}$ is the viewing direction, $t$ is the distance from the sampling point to the origin. In particular, the position coordinates and viewing direction need to be added to the encoding, and the encoding function is:
@@
The Encoding is useful for extracting high-dimensional information for feature enhancement and avoiding similarity.[29] Each pixel point radiates a ray and creates MLP (Multi-Layer Perception), which is trained and saves the weighting parameters. NeRF’s MLP has 10 layers, with 256 neurons in each of the first 9 layers and 128 neurons in the 10th layer. Then volume rendering is used to predict the colour values of the pixel points.[30] The volume rendering function is:
@@
where $T(t)$ denotes the transmittance function, which is affected by the opacity of the sampling points and the sampling distance:
@@
Since the sampled points are discrete, we need to take an integral of them. So we take an approximate integral to calculate the rendered color values.The approximate equation is:
@@
Finally, the loss is calculated from the rendered colour values and backpropagated:
@@
Neural Reconstruction by Neus: Neus is a reconstruction method that allows the surface of an object to be reconstructed more closely to the real surface. It takes the density values used by NeRF (Neural Radiance Fields) and replaces them with the point-to-surface distance (Signed Distance Field)[27]. The surface $\mathbf{S}$ of an object is the function equal to zero, the function is $\mathbf{S}=\left\{\mathrm{x}\in\mathbb{R}^{3}|f(\mathbf{x})=0\right\}.$ However, in 3D ray-tracing scenes, the use of the basic logistic density distribution $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}.$ introduces surface biases that affect the quality of the reconstruction. Therefore, Neus introduces an unbiased and occlusion-aware weight function:
@@
The Iterative Closest Point (ICP) algorithm iterates over the nearest points to obtain corresponding points. The point cloud data of different orientations and angles are unified under the same coordinate system through the Euler transformation relationship between corresponding points. This enriches the target point cloud data and makes it more complete. In 3D point cloud acquisition, the positional mapping between two sets of points is represented by displacement and rotation transformations totalling six degrees of freedom:
@@ III-D2 Measurement of Phenotypic data
The 3D model reconstructed using NeRF needs to be converted to actual size for final phenotype estimation. To achieve this, scaling markers that are regular in shape and relatively complete must be found and identified. Accurate calibration plates are used as markers for the calibration reconstruction. The scale factor between the estimated and actual lengths of the calibration plate reconstructed point cloud is then used as the corrected scale of the reconstruction. The estimation result is: